Johnson Company reports the following operating results for the month of August: Sales $441,000 (units 4,900), variable costs $261,000, and fixed costs $93,000. Management is considering the following independent courses of action to increase net income.

1. Increase selling price by 10% with no charge in total variable costs or units sold

2. Reduce variable costs to 58% of sales.

Compute the net income to be earned under each alternative.

1. Net income $____________

2. Net income $____________

Which course of action will produce the higher net income? Alternative 1 or alternative 2?
Answer #1

Sales = $441,000

Number of units sold = 4,900

Variable cost = $261,000

Fixed cost = $93,000

Selling price per unit = Sales/Number of units sold

= 441,000/4,900

= $90

Variable cost per unit = Variable cost/Number of units sold

= 261,000/4,900

= $53.27

1.

Increase in selling price = 10%

= 90 x 10%

= $9

Selling price per unit = 90 + 9

= $99

Profit = Sales - Variable cost - Fixed cost

= (4,900 x 99) - (4,900 x 53.27) - 93,000

= 485,100 - 261,023 - 93,000

= $131,077

2.

Variable cost = 58% of sales

= 90 x 58%

= $52.2

Profit = Sales - Variable cost - Fixed cost

= (4,900 x 90) - (4,900 x 52.2) - 93,000

= 441,000 - 255,780 - 93,000

= $92,220

First course of action would produce the higher net income.

Note : Exact answering figure may slightly differ due to rounding off.

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.